This question has been answered before but I need to be shown the steps in solving by using the ti84. An engineer is going to redesign an ejection seat for an airplane .The seat was designed for pilots weighing bet. 120-161 lbs. The new population of pilots has a normally distr. weights with a mean of 126 and SD of 33.1. Find the prob. they are bet. 120-161? I have tried using normalcdf but unsure what numbers to input where? We cannot use tables, so please show me as using the calculator. Please answer for statistics for Dummies!!
Answer by rothauserc(4718)

let X be the sample variable, then
:
Probability (P) ( 120 < X < 161 ) = P ( X < 161 ) - P ( X < 120 )
:
you need to calculate the z-score for each P
:
z-score(120) = (120 - 126) / 33.1 = -0.1813
z-score(161) = (161 - 126) / 33.1 = 1.0574
:
find the probability for each z-score, then subtract
:
Probability (P) ( 120 < X < 161 ) = 0.855 - 0.428 = 0.427
: